OpenStudy (anonymous):

Please help me with Ax + By = C standard form! Input in standard form the equation of the given line. The line with m = -3/2 and b = 6 Input in standard form the equation of the given line. The line through (0, -3) and (3, 0)

OpenStudy (anonymous):

Please help me!

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

Please help me!

OpenStudy (anonymous):

Okay, just making sure someone is really there.

OpenStudy (anonymous):

honestly this is pretty simple

OpenStudy (anonymous):

ok phew! I was worried I would never get help! So how do I start?

OpenStudy (anonymous):

do you know the point slope formula?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Do you mind writing it down/ posting it?

OpenStudy (anonymous):

Ok: (y1 - y2) / (x1 - x2)

OpenStudy (anonymous):

um not really, thats the formula to find a slope

OpenStudy (anonymous):

oh, what's the actual formula then?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

Does that look familiar?

OpenStudy (anonymous):

nevermind that doesn't matter

OpenStudy (anonymous):

It does like familiar though.

OpenStudy (anonymous):

so is there only one coordinate?

OpenStudy (anonymous):

I think so, this is all that I have to work with.

OpenStudy (anonymous):

(y-y1)=-3/2(x-6)

OpenStudy (anonymous):

Ok, what next?

OpenStudy (anonymous):

one sec, let me rearrange it.

OpenStudy (anonymous):

Sorry for making it more complicated than it should be. I just noticed that the first equation really is y=-3/2x+6

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

So now, I need to move the -3/2x over to the left.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It would be (3/2)x+y=6

OpenStudy (anonymous):

k, next problem?

OpenStudy (anonymous):

oh wait, I can't use fractions :(

OpenStudy (anonymous):

y=x-3 would be the equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If you can't use fractions, then just write 1.5x+y=6

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

what about on the second problem?

OpenStudy (anonymous):

1: 1.5x+y=6 2: y=x-3 ------> convert this into standard form

OpenStudy (anonymous):

ok so -x + y = 3?

OpenStudy (anonymous):

the second equation can be written as -x+y=-3 or x-y=3

OpenStudy (anonymous):

Okay, well, thanks :)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Do I get a medal?

OpenStudy (anonymous):

Oops sorry, I always give medals! :D